# Towards Adversarial Robustness via Debiased High-Confidence Logit Alignment

### v1

@@ Introduction
MART (Wang et al. 2020) 82.99 56.25 55.48 55.45 52.26 50.67 9.52
AWP (Wu, Xia, and Wang 2020) 82.67 57.80 57.21 57.07 54.82 51.90 6.90
FSR (Kim et al. 2023) 82.92 56.69 55.94 55.51 53.93 51.74 7.42
CFA (Wei et al. 2023) 84.43 57.87 56.90 56.64 54.60 51.85 10.36
UIAT (Dong et al. 2023) 82.94 58.66 58.12 58.05 54.11 52.17 7.92
SGLR (Li et al. 2024b) 85.76 57.53 56.91 56.66 54.28 52.07 9.38
@@
MART (Wang et al. 2020) 54.69 32.06 31.90 31.88 28.77 27.25 9.96
AWP (Wu, Xia, and Wang 2020) 57.94 34.01 33.75 33.72 30.74 28.90 7.87
FSR (Kim et al. 2023) 57.48 32.93 32.30 32.26 29.16 27.04 7.84
CFA (Wei et al. 2023) 60.92 33.10 32.56 32.41 30.49 28.04 10.47
UIAT (Dong et al. 2023) 57.65 34.27 33.91 33.85 30.97 29.03 11.70
SGLR (Li et al. 2024b) 61.02 33.43 32.98 32.82 30.72 28.50 15.67
@@
[…]
MART (Wang et al. 2020) 81.17 59.71 59.23 59.03 57.17 56.18 15.11
AWP (Wu, Xia, and Wang 2020) 83.25 61.32 60.92 60.74 59.39 57.05 25.65
FSR (Kim et al. 2023) 85.86 60.95 60.31 59.70 57.86 56.31 21.95
CFA (Wei et al. 2023) 83.64 60.59 60.18 60.03 58.26 56.70 23.78
UIAT (Dong et al. 2023) 81.35 62.49 62.54 62.51 61.10 57.84 17.87
SGLR (Li et al. 2024b) 84.99 61.48 60.90 60.64 58.39 57.11 22.05
DHAT (Ours) 82.11 64.06 63.72 63.55 62.63 59.09 12.63
DHAT-AWP (Ours) 83.78 64.57 64.34 64.20 63.45 59.47 16.77
DHAT-CFA (Ours) 84.28 64.91 64.92 64.77 63.58 59.95 15.76
@@
Experimental Result
Robust performance.
DHAT consistently demonstrates superior robustness across all datasets. For example, on CIFAR-10, DHAT outperforms the second-best method by 1.93% and 1.16% under PGD-10 and C&W attacks, respectively. DHAT also shows the best performance on more complex datasets such as CIFAR-100 and high-resolution Imagenette datasets. Specifically, on Imagenette, DHAT achieves improvements of 1.57% and 1.25% over the second-best method under PGD-10 and AA attacks, respectively. Transfer attack performance comparisons are also reported in the supplementary material.
Combine with other AT methods.
Robust overfitting.
Generalization Analysis
Performance with diverse network architecture.
@@
MART (Wang et al. 2020) 75.05 49.05 4.57 66.72 44.27 3.60 76.05 49.59 5.91
AWP (Wu, Xia, and Wang 2020) 75.59 50.80 3.91 67.14 45.97 3.47 76.74 50.46 4.07
FSR (Kim et al. 2023) 78.38 50.59 3.97 69.59 44.90 3.30 77.28 50.02 4.73
CFA (Wei et al. 2023) 77.88 50.64 4.27 70.45 45.53 4.66 76.70 49.97 4.46
UIAT (Dong et al. 2023) 78.42 51.00 2.99 68.46 45.27 3.57 75.19 51.23 3.70
SGLR (Li et al. 2024b) 80.39 50.55 5.21 70.90 45.04 3.94 80.62 50.73 5.00
@@
Different attack magnitude.
Figure 3 illustrates the performance of DHAT under varying magnitudes of PGD-10 and C&W attacks. Our findings reveal that DHAT maintains a significant performance advantage across all tested attack magnitudes. Compared to other methods, DHAT exhibits improved stability and robustness, with minimal performance degradation even under more severe attacks. This consistent performance across varying attack levels underscores the effectiveness of our approach in mitigating the impact of adversarial perturbations. Further details of the attack analysis are provided in the supplementary material.
Ablation Analysis
@@
Impact of individual components.
@@
Analysis of parameter sensitivity.
@@
Analysis of threshold factor for quantifying spurious bias.
@@
Figure 6: Visualization of feature activation maps using UIAT and our proposed method DHAT, utilizing Grad-CAM under inverse attack ($\epsilon=-8$) and adversarial attack ($\epsilon=8$). Left-to-right: the original image; normal feature activation maps; feature activation maps under adversarial attacks for UIAT and DHAT; and feature activation maps under inverse adversarial attacks for UIAT and DHAT.
Discussion DHAT Effectiveness
@@
We begin by visualizing the feature activation maps of models trained with UIAT and DHAT under both adversarial attacks and inverse adversarial attacks. These visualizations, as depicted in Figure 6, highlight notable differences in feature attention maps between the models. Specifically, our proposed method DHAT restores the model’s attention to its normal state. In contrast, UIAT exhibits biased feature activation, particularly under inverse adversarial attack conditions, where the model’s attention shifts toward background features, reducing its attention to foreground features. DHAT’s effectiveness is attributed to its ability to recalibrate biased high-confidence logits during inverse adversarial attacks and reduce the correlation between high-confidence logits and background features in affine space. This ensures that the alignment with inverse adversarial examples is not compromised by spurious correlation bias.
@@
(a) IoU with True Foreground (b) IoU with True Background
Figure 7: Intersection over Union (IoU) of attention regions with true foreground and background regions. The comparison is conducted among standard adversarial training (AT), UIAT, and DHAT.
Furthermore, we quantitatively evaluate DHAT’s impact on mitigating spurious correlation bias, we compute the Intersection over Union (IoU) between the model’s attention maps and the true foreground and background regions. Figure 7 illustrates the comparison under inverse adversarial attacks using conventional AT, UIAT, and DHAT models. Our results show that DHAT substantially improves the model’s attention to foreground features while effectively alleviating biased feature activation. Specifically, DHAT achieves higher IoU scores with true foreground regions and lower IoU scores with background regions compared to both AT and UIAT. These findings confirm that DHAT effectively mitigates spurious correlation bias.
Conclusion
In this work, we identify a critical issue in adversarial training: AT model exhibits biased feature activation under inverse adversarial attacks. Specifically, training with inverse adversarial examples leads to a shift in the model’s attention toward background features, resulting in spurious correlation bias. To tackle this challenge, we propose a novel method called Debiased High-Confidence Adversarial Training (DHAT). DHAT incorporates two pivotal techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and recalibrates biased high-confidence logits to align the logits of adversarial examples with these debiased high-confidence logits. FLOE further restores the model’s attention to its normal state by reducing the correlation between high-confidence logits and background features in affine space. This work highlights DHAT’s effectiveness in enhancing model resilience and presents an adversarial training framework that integrates seamlessly with advanced training strategies.
Supplementary
This supplementary material provides an in-depth examination of the methodologies and results that support the experiments detailed in our study. It aims to offer comprehensive information that validates the findings and ensures reproducibility and transparency.
Experimentation Details
@@ Conclusion
Datasets Information
CIFAR-10 (Krizhevsky, Hinton et al. 2009) comprises 10 classes, with each image having a resolution of 32x32 pixels. This dataset is commonly used for basic image classification tasks and serves as a benchmark for evaluating the performance of various models on a relatively simple dataset.
CIFAR-100 (Krizhevsky, Hinton et al. 2009) contains 100 classes, with each image also having a resolution of 32x32 pixels. The expanded class set in CIFAR-100 increases complexity, enabling a more rigorous evaluation of models in distinguishing among a broader range of objects.
Imagenette (Howard 2019), a subset of ImageNet, includes 10 classes with images at a higher resolution of 320x320 pixels. This dataset is constructed to provide a more realistic and challenging benchmark due to its increased resolution and closer representation of real-world images.
Assessing models on datasets with a larger number of classes and higher resolutions, such as CIFAR-100 and Imagenette, facilitates in-depth analysis of their generalization capabilities. This ensures that models demonstrate effectiveness not only on simpler tasks but also exhibit resilience when confronted with more complex and diverse data.
Training Setup
In the training setup, we employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$. The initial learning rate is set to 0.1 for CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009) datasets, while 0.01 for Imagenette (Howard 2019). All models undergo 100 epochs of training, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs. The trade-off hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 and 1.0. To maintain experimental fairness, all comparative methods undergo identical training strategy, involving the same number of epochs and the same optimizer. For Grad-CAM to quantify the spurious correlation bias, we use a pre-trained ResNet-18 model under simple training, introducing minimal additional training cost. The experiments were conducted on a single NVIDIA Tesla A100.
@@
For adversarial training, we create adversarial examples through a 10-iteration attack where the maximum $\ell_{\infty}$-norm of the adversarial perturbation is limited to $\epsilon=8/255$, using a step size of $\alpha=2/255$. To ensure the reliability of our experiments, we align the inverse adversarial example generation process with UIAT settings (Dong et al. 2023), employing the same loss function and same perturbation constrained by $\epsilon=4/255$.
@@
Notably, our method does not employ label smoothing techniques. This choice is deliberate to avoid potential conflicts with efforts to mitigate spurious correlation bias, as label smoothing can blur class boundaries. Additionally, unlike UIAT, our approach does not utilize momentum terms to stabilize the generation process of inverse adversarial examples, focusing solely on the intrinsic properties of our method.

### v2

@@ 4 Experiments
MART [45] 62.31 43.56 42.89 42.54 40.62 38.94 13.76
AWP [49] 64.25 45.13 44.67 44.21 41.96 40.02 12.82
FSR [21] 64.70 44.31 43.82 43.45 41.23 39.30 14.17
CFA [47] 65.38 44.87 44.32 43.96 41.58 39.65 15.49
UIAT [9] 62.64 45.29 44.85 44.57 42.13 40.18 14.68
SGLR [27] 63.46 44.52 44.05 43.78 41.36 39.47 19.72
DHAT (Ours) 65.90 46.83 46.42 46.11 43.25 41.70 9.53
DHAT-AWP (Ours) 65.62 47.38 47.04 46.79 43.62 42.07 10.27
DHAT-CFA (Ours) 66.26 48.27 47.89 47.53 44.18 42.45 11.64
[…]
MART [45] 82.99 56.25 55.48 55.45 52.26 50.67 9.52
AWP [49] 82.67 57.80 57.21 57.07 54.82 51.90 6.90
FSR [21] 82.92 56.69 55.94 55.51 53.93 51.74 7.42
CFA [47] 84.43 57.87 56.90 56.64 54.60 51.85 10.36
UIAT [9] 82.94 58.66 58.12 58.05 54.11 52.17 7.92
SGLR [27] 85.76 57.53 56.91 56.66 54.28 52.07 9.38
@@
MART [45] 54.69 32.06 31.90 31.88 28.77 27.25 9.96
AWP [49] 57.94 34.01 33.75 33.72 30.74 28.90 7.87
FSR [21] 57.48 32.93 32.30 32.26 29.16 27.04 7.84
CFA [47] 60.92 33.10 32.56 32.41 30.49 28.04 10.47
UIAT [9] 57.65 34.27 33.91 33.85 30.97 29.03 11.70
SGLR [27] 61.02 33.43 32.98 32.82 30.72 28.50 15.67
@@
4.1 Experimental Settings
4.1.1 Datasets and models.
We conduct experiments on the CIFAR-10, CIFAR-100 [23], and ImageNet-1K [31] datasets. The architecture for our experiments include the mainstream models WRN28-10 [52], ResNet-18 [16], ResNet-50 [16], Inception-V3 [43], and VGG16 [38]. Detailed information is provided in the supplementary material.
4.1.2 Training setup
For CIFAR datasets, each model is trained for 100 epochs, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs. For adversarial example generation, we use a fixed 10-iteration attack. The maximum $\ell_{\infty}$-norm of the adversarial perturbation is set to $\epsilon=8/255$, while the inverse adversarial perturbation is set to $\epsilon=4/255$. Both perturbations use a fixed step size of $\alpha=2/255$. For the ImageNet dataset, we adopt the approach in [39] with 2-iteration PGD over 50 epochs, and set $\epsilon=4/255$ and $\alpha=1/255$ for adversarial perturbations. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 in all data sets. Comprehensive training details are available in the supplementary material.
4.1.3 Evaluation setup.
For robustness performance evaluation, we employ PGD [29], C&W [4], and AutoAttack (AA) [7] attacks under the $\ell_{\infty}$ norm. AutoAttack comprises APGD-DLR [7], APGD-CE [7], FAB [8], and Square [2]. “Clean” denotes natural examples without adversarial perturbations. Additionally, we introduce the concept of the robust generalization gap, referred to as the “Robust Gap”, which measures the difference in robustness performance between the training and test sets as a direct indicator of the generalization ability [51]. This metric quantifies the strength of spurious correlations, where larger gaps indicate greater reliance on features that lack causal relationships with target labels.
@@ 4.1.2 Training setup
MART [45] 75.05 49.05 4.57 66.72 44.27 3.60 76.05 49.59 5.91
AWP [49] 75.59 50.80 3.91 67.14 45.97 3.47 76.74 50.46 4.07
FSR [21] 78.38 50.59 3.97 69.59 44.90 3.30 77.28 50.02 4.73
CFA [47] 77.88 50.64 4.27 70.45 45.53 4.66 76.70 49.97 4.46
UIAT [9] 78.42 51.00 2.99 68.46 45.27 3.57 75.19 51.23 3.70
SGLR [27] 80.39 50.55 5.21 70.90 45.04 3.94 80.62 50.73 5.00
@@
4.2 Experimental Result
4.2.1 Robust performance.
DHAT consistently demonstrates superior robustness across all datasets. For example, on CIFAR-10, DHAT outperforms the second-best method by 1.93% and 1.16% under PGD-10 and C&W attacks, respectively. DHAT also shows the best performance on more complex datasets such as CIFAR-100 and high-resolution Imagenette datasets. Specifically, on ImageNet-1K, DHAT achieves improvements of 1.54% and 1.52% over the second-best method under PGD-10 and AA attacks, respectively.
4.2.2 Combine with other AT methods.
4.2.3 Robust overfitting.
4.3 Generalization Analysis
4.3.1 Performance with diverse network architecture.
@@ 4.3 Generalization Analysis
4.3.2 Different attack magnitude.
Figure 3 illustrates the performance of DHAT under varying magnitudes of PGD-10 and C&W attacks. Our findings reveal that DHAT maintains a significant performance advantage across attack magnitudes. Compared to other methods, DHAT exhibits improved stability and robustness, with minimal performance degradation even under more severe attacks. This consistent performance across varying attack levels underscores the effectiveness of our approach in mitigating the impact of adversarial perturbations. Further details are provided in the supplementary material (Section 7).
Table 3: Comparison of computational costs (# A100 GPU hours) using different adversarial training methods on CIFAR-10.
Method Architecture
ResNet-18 WRN28-10 Inception-V3
MART [45] 0.050 0.285 0.331
AWP [49] 0.048 0.274 0.318
FSR [21] 0.051 0.291 0.337
CFA [47] 0.045 0.262 0.303
UIAT [9] 0.040 0.245 0.283
SGLR [27] 0.039 0.240 0.278
DHAT (Ours) 0.041 0.248 0.287
4.4 Computational Efficiency
The computational costs of our method and baseline models are provided in terms of per-epoch training time on a single NVIDIA A100 GPU with a batch size of 128 on the CIFAR-10 dataset, as shown in Table 3. Our method demonstrates significantly improved time efficiency compared to the baseline MART, AWP, FSR, and CFA. While it requires slightly more computational resources than the state-of-the-art methods UIAT and SGLR, this minor increase is justifiable given the performance gains achieved.
4.5 Ablation Analysis
@@
4.5.1 Impact of individual components.
@@
4.5.2 Analysis of parameter sensitivity.
@@
4.5.3 Analysis of threshold factor for quantifying spurious bias.
@@
Figure 6: Visualization of feature activation maps using UIAT and DHAT, utilizing Grad-CAM under inverse attack ($\epsilon=-8$) and adversarial attack ($\epsilon=8$). Left-to-right: the original image; normal feature activation maps; feature activation maps under adversarial attacks for UIAT and DHAT; and feature activation maps under inverse adversarial attacks for UIAT and DHAT.
4.6 Discussion DHAT Effectiveness
@@
4.6.1 Feature Activation Map Analysis
Figure 6 visualizes feature activation maps of UIAT and DHAT models under adversarial and inverse adversarial attacks. DHAT successfully maintains normal attention patterns, while UIAT exhibits biased activation with attention shifting significantly to background features during inverse attacks. DHAT achieves this robust performance by recalibrating biased high-confidence logits and systematically reducing correlation between logits and background features.
@@ 4.6.1 Feature Activation Map Analysis
(a) IoU with Normal Foreground (b) IoU with Normal Background
Figure 7: IoU of attention regions with normal foreground and background regions. The comparison is conducted among conventional AT, UIAT, and DHAT on the Imagenette Dataset.
Table 5: Comparison of training with different magnitude reverse adversarial samples (UIAT) and our debiased training DHAT using WRN28-10 on the CIFAR-10.
[…]
UIAT -4 90.53 61.45 82.94 52.07 9.38
UIAT -8 92.57 62.87 82.26 51.33 11.54
UIAT -12 94.90 65.28 81.72 50.85 14.43
DHAT (ours) -4 88.73 56.61 83.95 53.10 3.51
4.6.2 Quantitative Evaluation of Bias Mitigation
Figure 7 quantifies DHAT’s effectiveness through IoU between attention maps and true foreground/background regions. DHAT achieves higher foreground IoU and lower background IoU than both AT and UIAT, confirming its superior feature attention. Table 5 shows stronger inverse attacks in UIAT worsen overfitting, increasing the robust gap from 9.38 to 14.43. In contrast, DHAT achieves better test robustness (83.95% PGD-10, 53.10% AA) with a significantly smaller robust gap (3.51), effectively mitigating overfitting while maintaining strong performance.
5 Conclusion
In this work, we identify a critical issue in adversarial training: AT model exhibits biased feature activation under inverse adversarial attacks. Specifically, training with inverse adversarial examples causes the model attention to shift from the foreground to background features, resulting in spurious correlation bias. To tackle this challenge, we propose a novel method called Debiased High-Confidence Adversarial Training (DHAT). DHAT incorporates two pivotal techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and recalibrates biased high-confidence logits to align the logits of adversarial examples with these debiased high-confidence logits. FLOE further restores the model’s attention to its normal state by reducing the correlation between high-confidence logits and background features in affine space. This work highlights DHAT’s effectiveness in enhancing model resilience and presents an adversarial training framework that integrates seamlessly with advanced training strategies.
Acknowledgment
This work is supported by the National Natural Science Foundation of China under Grant No. 62276221, 62376232; Key Laboratory of Equipment Data Security and Guarantee Technology, Ministry of Education under Grant No. 2024020200.
Supplementary Materials
This supplementary material offers a detailed examination of the methodologies and results that underpin the experiments in our study. It is designed to provide comprehensive information to validate the findings and ensure reproducibility and transparency.
The content is organized as follows:
  • Section 6: A comprehensive and detailed description of the experimental setup and methodology.
  • Section 7: A formal presentation of the proposed algorithm, including its pseudocode representation.
  • Section 8: In-depth evaluation of model performance under various transfer attack scenarios.
  • Section 9: Detailed results from adversarial training using synthetic data for robustness assessment.
  • Section 10: Extensive exploration of techniques for generating foreground-background attention maps.
  • Section 11: Thorough examination of the model’s robustness under varying levels of attack intensity.
6 Experimentation Details
@@
6.1 Training Setup
6.1.1 CIFAR-10 and CIFAR-100
For the CIFAR-10 and CIFAR-100 [23] datasets, we employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$. The initial learning rate is set to 0.1, and models undergo 100 epochs of training with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs.
@@ 6.1.1 CIFAR-10 and CIFAR-100
For adversarial training on these datasets, we create adversarial examples through a 10-iteration attack where the maximum $\ell_{\infty}$-norm of the adversarial perturbation is limited to $\epsilon=8/255$, using a step size of $\alpha=2/255$. To ensure reliability and fairness, we align the inverse adversarial example generation process with the UIAT settings [9], employing the same loss function and perturbation constrained by $\epsilon=4/255$.
6.1.2 ImageNet-1K
For the ImageNet-1K dataset, we strictly follow the training protocol established in [39] to ensure fair comparison with existing methods. Specifically, we implement a 2-iteration PGD over 50 epochs, and set $\epsilon=4/255$ and $\alpha=1/255$ for adversarial perturbations. This adherence to established benchmarking parameters enables direct and meaningful comparison with state-of-the-art approaches evaluated on this challenging large-scale dataset.
6.1.3 Common Settings
The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are consistently set to 1.0 across all datasets. To maintain experimental fairness, all comparative methods undergo identical training strategies with the same number of epochs and optimizer configurations. For Grad-CAM to quantify the spurious correlation bias, we use a pre-trained ResNet-18 model under simple training, introducing minimal additional training cost.
@@ 6.1.3 Common Settings
All experiments were conducted on a single NVIDIA Tesla A100. Notably, our method does not employ label smoothing techniques to avoid potential conflicts with efforts to mitigate spurious correlation bias, as label smoothing can blur class boundaries and undermine the precision required for robust decision-making. Additionally, unlike UIAT, our approach does not utilize momentum terms to stabilize the generation process of inverse adversarial examples, instead focusing solely on the intrinsic properties and underlying dynamics of our method. Comprehensive training details are available in the supplementary material.
6.2 Evaluation Setup
Our evaluation consists of two main components: robustness performance and robust generalization. To evaluate robustness performance, we use PGD [29], C&W [4], and AutoAttack (AA) [7] within the $\ell_{\infty}$-norm. AutoAttack includes several attack methods such as APGD-DLR [7], APGD-CE [7], FAB [8], and Square [2]. Adversarial attacks are generated using a step size of $\alpha=2/255$ and a specified maximum $\ell_{\infty}$-norm. Note that “Clean” indicates natural examples unaffected by adversarial perturbations.
To evaluate robust generalization, we introduce the concept of the robust generalization gap [51] denoted as the “Robust Gap” quantifying the difference in robust performance between training and test sets under adversarial attacks. A smaller robust gap indicates improved robust generalization, reflecting reduced vulnerability to robust overfitting in the adversarial training model.
7 Algorithm Details
Algorithm 1 Debiased High-Confidence Adversarial Training (DHAT)
0:  Training dataset $(X,Y)=\{(x_{i},y_{i})\}_{i=1}^{N}$; Network parameters $\theta$; Perturbation budget $\epsilon$; Hyperparameters $\lambda_{1},\lambda_{2}$; Attention threshold $\omega$.
0:  Robust model parameters $\theta^{*}$
1:  Initialize network parameters $\theta$
2:  for each epoch do
3:     for each mini-batch $(x,y)$ do
4:        // Generate adversarial examples
5:        $\hat{x}\leftarrow\arg\max_{||x^{\prime}-x||_{p}\leq\epsilon}\mathcal{L}_{AT}(f_{\theta}(x^{\prime}),y)$
6:        $\hat{z}\leftarrow f_{\theta}(\hat{x})$
7:        // Generate inverse adversarial examples
8:        $\check{x}\leftarrow\arg\min_{||x^{\prime}-x||_{p}\leq\epsilon}\mathcal{L}_{Inv}(f_{\theta}(x^{\prime}),y)$
9:        $\check{z}\leftarrow f_{\theta}(\check{x})$
[…]
11:        $M\leftarrow\mathcal{A}(x)$Different attention map computation techniques}
12:        // Extract background features from inverse adversarial examples
13:        $[\check{x}_{(B)}]_{(i,j)}\leftarrow\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)}$
14:        $\check{z}_{(B)}\leftarrow f_{\theta}(\check{x}_{(B)})$
15:        // Compute debiased high-confidence logits
16:        $\check{z}^{*}\leftarrow\check{z}-\check{z}_{(B)}$
17:        // Compute DHLR loss
18:        $\mathcal{L}_{DHLR}\leftarrow\mathcal{L}_{KL}(\phi(\check{z}^{*})||\phi(\hat{z}))$
19:        // Compute FLOE loss
20:        $\mathcal{L}_{FLOE}\leftarrow-|\check{z}-\frac{\check{z}\cdot\check{z}_{(B)}}{|\check{z}_{(B)}|^{2}}\cdot\check{z}_{(B)}|_{p}$
21:        // Compute total loss
22:        $\mathcal{L}_{DHAT}\leftarrow\mathcal{L}_{AT}(\hat{z},y)+\lambda_{1}\cdot\mathcal{L}_{DHLR}+\lambda_{2}\cdot\mathcal{L}_{FLOE}$
23:        // Update network parameters
24:        $\theta^{*}\leftarrow\theta-\nabla_{\theta}\mathcal{L}_{DHAT}$
25:     end for
26:  end for
27:  return  $\theta^{*}$
Algorithm 1 outlines the complete workflow of our proposed Debiased High-Confidence Adversarial Training (DHAT) framework, which systematically mitigates spurious correlations in adversarial training through two primary components: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE).
To efficiently compute attention maps, we employ Grad-CAM (Line 11) due to its computational efficiency. However, our framework is compatible with alternative saliency-based methods, such as SAM [22], which can provide enhanced performance at a higher computational cost. The extraction of background features (Lines 13-14) is performed using an adaptive threshold $\omega$, ensuring the identification of non-discriminative regions that contribute to spurious correlations. The overall training objective (Line 23) integrates standard adversarial training with our proposed debiasing regularization terms, weighted by $\lambda_{1}$ and $\lambda_{2}$.
Despite incorporating additional minimal computational steps, the DHAT framework introduces only a marginal increase in resource consumption compared to standard adversarial training while significantly enhancing model robustness. Crucially, DHAT achieves these improvements without compromising clean-data accuracy, making it a highly practical and scalable solution for real-world adversarial defense applications.